Baddie156:

Simplify the expression: 8 ( 5w + 7) + 9 (4w - 7)

rocfp:

1.Remove any grouping symbol such as brackets and parentheses by multiplying factors. 2.Use the exponent rule to remove grouping if the terms are containing exponents. 3.Combine the like terms by addition or subtraction. 4. Combine the constants

Baddie156:

ok

RandosWorld:

First you remove the parentheses.... Then you collect like terms and calculate and then you get your answer. I can go into more depth if you would like.

supie:

\(8 ( 5w + 7) + 9 (4w - 7)\) The first thing we do is distribute \(8 ( 5w + 7) \&\ 9 (4w - 7)\) We distribute what is outside the parenthesis to everything that is inside the parenthesis so for \(8 ( 5w + 7)\) we do \(8 \times 5w= 40w\) and \(8 \times 7= 56\) so \(8 ( 5w + 7)\) becomes \(45w+56\) and for \(9(4w-7)\) is \(9 \times 4w = 36w\) and \(9 \times 7 =63\) So \(9 (4w - 7)\) turns into \(36w+63\) so now we have \(40w+56+36w+−63\) now combine like terms \((40w+36w)+(56+−63)=?\)

Baddie156:

@supie thx

supie:

np

Baddie156:

@supie really helpful 🥰